# To reset, or not to reset – that is the question

### v1

@@ Abstract
Whether to reset qubits, or not, during quantum error correction experiments is a question of both foundational and practical importance for quantum computing. Text-book quantum error correction demands that qubits are reset after measurement. However, fast qubit reset has proven challenging to execute at high fidelity. Consequently, many cutting-edge quantum error correction experiments are opting for the no-reset approach, where physical reset is not performed. It has recently been postulated that no-reset is functionally equivalent to reset procedures, as well as being faster and easier. For memory experiments, we confirm numerically that resetting provides no benefit. On the other hand, we identify a remarkable difference during logical operations. We find that unconditionally resetting qubits can reduce the duration of fault-tolerant logical operation by up to a factor of two as the number of measurement errors that can be tolerated is doubled. We support this with numerical simulations. However, our simulations also reveal that the no-reset performance is superior if the reset duration or infidelity exceeds a given threshold. Lastly, we introduce two novel syndrome extraction circuits that can reduce the time overhead of no-reset approaches. Our findings provide guidance on how experimentalists should design future experiments.
@@
The structure of our paper is as follows. In Section I, we describe how the decoding problem changes when we switch from the unconditional-reset to the conditional-reset or no-reset scheme. In Section II, we first describe a Pauli noise model, inspired by currently available realistic superconducting hardware, that we use throughout (Section II.1). Then, we present numerical results that verify the negligible difference between the three reset schemes for quantum memory with the rotated planar code (Section II.2). After that, we describe the stability experiment and why it is a useful proxy for estimating the time cost of FTQC. Furthermore, we demonstrate through numerical simulations with the stability experiment that FTQC incurs a substantial time overhead in the no-reset (and conditional-reset) scheme, provided resetting is fast enough and is of high enough fidelity (Section II.3). In Sections III and IV, we present two alternative syndrome extraction circuits in the no-reset scheme that overcome this time overhead problem. Even though we present circuits only for the planar code, they can also be applied straightforwardly to more general QEC codes. With Section V, we conclude the paper by discussing the obtained results and possible future research directions.
@@ I The structure of detectors in the three reset schemes
Now, we analyse which detectors are triggered (change value from the noise-free case) by different errors. A Pauli error on a data qubit occurring during the initialisation of the auxiliary qubits triggers only detectors with the same time coordinate. We refer to errors with this property as “space-like”. On the other hand, a (classical or quantum) measurement error on the auxiliary qubit of $g$ in the $j$th QEC round triggers at most two detectors, namely at coordinates $(x,y,j-1)$ and/or $(x,y,j)$. Such an error is described as “time-like”. Some Pauli errors that occur mid-syndrome-extraction-circuit trigger a set of detectors where neither the spatial nor the time coordinates coincide for all the detectors – these are known as hook errors [24].
@@
This structural difference of detectors does not change the triggered detectors corresponding to any mid-circuit quantum Pauli error. However, misclassification of a measurement outcome results in a different combination of triggered detectors to the unconditional-reset case. To see this, let us assume we have a classification error on measurement result $n_{g,j}$, i.e., we read out $n_{g,j}\oplus 1$, even though the qubit collapsed into the $n_{g,j}$-eigenstate. This triggers two detectors at coordinates $(x,y,j-1)$ and $(x,y,j+1)$ (cf. Equation 2), so in the decoding graph it corresponds to a time-like edge with length $2$.
In Figure 2, we present a topological perspective on how vertical strings of classification errors can lead to an undetectable logical failure in lattice surgery and stability experiments, analogous to how horizontal strings of physical qubit errors do the same in quantum memory experiments. Considering these vertical failure mechanisms, we see that the no-reset approach will tolerate only half as many classification errors as the unconditional-reset approach, since each error has twice the vertical length in the former scheme. This is not the case for quantum memory; however, as we will see in Section II.3, this difference significantly impacts the performance of the stability experiment, and hence FTQC.
@@ II.1 Noise model
We assume that the device has fixed and uniform qubit connectivity, with a planar nearest-neighbour architecture connecting each (bulk) qubit to four others. This so-called square-grid connectivity enables implementation of the planar code and is found in several modern superconducting quantum computers [e.g., 27, 14]. In Section IV, we will present a circuit that requires slightly different planar connectivity – some qubits need only be connected to three others instead of four.
@@
We construct our simulated circuits [28] in terms of the controlled-$Z$ two-qubit gate (CZ) and single-qubit $X$- and $Z$-basis rotations, a common native gate set for superconducting qubits [e.g., 27]. In particular, we use the $\pi/2$ rotations $\sqrt{X}$ and $S$, given by
@@
$T_{1}=\frac{p}{p^{\mathrm{ref}}}T_{1}^{\mathrm{ref}},\quad T_{2}=\frac{p}{p^{\mathrm{ref}}}T_{2}^{\mathrm{ref}}$ (6)
@@ II.2 The impact of no-reset on memory
Figure 4: Comparison of quantum memory performances in the unconditional-reset (dots and solid lines) and no-reset (triangles and dashed lines) schemes using standard syndrome extraction circuits. For each distance $d=5,7,9$ (corresponding to red, blue, black colours), we performed $d$ QEC rounds. As is visible, there is negligible difference between the two schemes, confirming the statement made in [13, Sec. S2].
Using the noise model described in Section II.1, we performed quantum memory simulations for distance-$d$ rotated planar code for $d$ number of QEC rounds. Figure 4 compares the unconditional-reset and no-reset schemes. We constructed circuits using the python library stim [31] and took samples which were then decoded using the MWPM python library pymatching [25]. The circuits are available here: [28]. The number of samples for each data point was the minimum required to observe $10^{4}$ logical failures or reach $10^{8}$ samples, whichever happened first. This ensures error bars that are too small to be visible in our plots. We calculated the sample logical failure probabilities $p_{X}$ and $p_{Z}$, for $X$ and $Z$-memory, respectively, and combined them into one quantity as $p_{L}=p_{X}+p_{Z}-p_{X}p_{Z}$. Our results (Figure 4) confirm the statement of [13], mentioned earlier, that there is negligible difference between the two reset schemes in the context of quantum memory. More precisely, the no-reset scheme’s performance is slightly better, as the threshold is somewhat higher and the logical failure probabilities are slightly lower.
@@
Figure 5: Comparison of the stability experiment performances in the two reset schemes using standard syndrome extraction circuits. Colours indicate different physical error probabilities: $p=10^{-2}$ (blue), $p=10^{-2.5}$ (orange), $p=10^{-3}$ (purple). (a) Logical failure probabilities of the stability experiment for the $4\times 4$ patch shown in Figure 1(b) plotted against the number of QEC rounds. Error bars corresponding to 3×3$\times$3 × the standard error of the mean are also shown as shaded areas. Note that for the $p=10^{-2}$ case the plots overlap. The unconditional reset is plotted with dots, the no-reset with triangles. Best $\log$-line fits are also shown with solid and dashed lines, respectively. (b) The time overhead $R_{\mathrm{ur}}$ that the unconditional-reset scheme requires when compared to the no-reset scheme given $t_{\mathrm{res}}=0$ (instantaneous), $100$ (fast) and $500$ ns (slow) resets, respectively. In particular, the first column of dots in the first plot corresponds to (a), since instantaneous reset implies equal execution time in the two reset schemes. The pink dashed line shows the break-even point, i.e. when $R_{\mathrm{ur}}=1$, below which the unconditional reset has better FTQC performance.
@@ II.3 The impact of no-reset on FTQC
Based on this, we would expect the stability performance to be improved by using unconditional resets – assuming fast enough and high-enough fidelity reset gates. We performed simulations to assess this. We considered planar code stability patches of sizes $w\times w$ with $w\in\{4,6,8\}$ and physical error rates $p\in\{10^{-2},10^{-2.5},10^{-3}\}$. For each, we prepared stim circuits [28], sampled, and decoded until we reached either $10^{9}$ total shots or $10^{3}$ logical failures, whichever happened first. The obtained logical failure probabilities plotted against the number of QEC rounds, $n=5,7,9,11,13$, are shown in Figure 5(a) for the $w=4$ case. They clearly show that implementing unconditional reset improves the performance for $p=10^{-2.5},10^{-3}$. Furthermore, this improvement increases as $p$ decreases.
@@
where $a$ and $\gamma$ both depend on the width of the patch $w$ [21]. Note that one QEC round takes $t_{\mathrm{nr}}=840$ ns with the no-reset scheme and, if the reset duration is $t_{\mathrm{res}}$ ns, one QEC round takes $t_{\mathrm{ur}}=(840+t_{\mathrm{res}})$ ns with the unconditional-reset scheme. These two times are equal only with instantaneous reset, i.e., $t_{\mathrm{res}}=0$ ns. Therefore, as a fairer comparison, Equation 7 may be expressed in terms of the time $t$ the experiment requires, instead of the number of QEC rounds, so that
@@
where subscripts $\mathrm{nr}$ and $\mathrm{ur}$ indicate the no- and unconditional-reset schemes, respectively. We define the time overhead of implementing unconditional resets as $R_{\mathrm{ur}}=\gamma_{\mathrm{nr}}/\gamma_{\mathrm{ur}}$ [32, Sec. 5.2]. This can be interpreted as follows: if we perform a stability experiment for $t_{\mathrm{nr}}$ ns without unconditional resets then, in order to match this performance using unconditional resets, we need to spend approximately $t_{\mathrm{ur}}=R_{\mathrm{ur}}t_{\mathrm{nr}}$ ns. If $R_{\mathrm{ur}}<1$, then implementing unconditional resets decreases the time cost of FTQC, and hence is worth implementing; otherwise not. The time overhead is plotted in Figure 5(b), assuming instantaneous reset $t_{\mathrm{res}}=0$ ns, fast reset $t_{\mathrm{res}}=100$ ns, and slow reset $t_{\mathrm{res}}=500$ ns. We call $p_{\mathrm{br}}$ the break-even point, where $R_{\mathrm{ur}}=1$. Clearly, as $t_{\mathrm{res}}$ increases, $p_{\mathrm{br}}$ decreases. Furthermore, it can be seen in Figure 5(b) that, as $p$ decreases, the time overhead decreases too, indicating that when $p$ is small enough, implementing unconditional reset improves FTQC performance. In all cases of instantaneous and fast reset, the break-even point satisfies $10^{-2.5}<p_{\mathrm{br}}<10^{-2}$. With slow reset, the break-even point is not visible in our plot, and we expect it to be at very low $p$. This indicates that, unless the reset gate error rate and duration are below certain values, implementing unconditional reset may not be beneficial for FTQC. However, if the break-even point is reached, either with decreased reset duration or higher-fidelity reset gates, the unconditional-reset scheme substantially improves FTQC performance.
@@ III Recovering time-like distance by spreading classification errors
The idea is to spread the auxiliary qubit measurement classification errors to the data qubits, thereby triggering additional detectors and thus requiring more errors for an undetectable time-like logical failure. We achieve this by applying two conditional Pauli gates immediately after measuring the auxiliary qubit – one conditioned on the measurement outcome and the other conditioned on the auxiliary qubit state itself; see Figure 6 for an $X$-stabiliser circuit. Note that the controlled gates are of $Z$-($X$-)type for $X$-($Z$-)type stabilisers, and that the classically-controlled gates do not have to be applied on the device – instead, their effect can be tracked in software. In the absence of a measurement classification error, the two conditional gates cancel so there is no overall effect on the data qubits. However, if there is a classification error, exactly one of the gates will have an effect, resulting in a Pauli gate being applied to the data qubit which then behaves like a Pauli error on the qubit. Therefore, misclassification of measurement $m_{g,j}$ will now trigger four detectors – the two which directly result from the measurement classification error, at coordinates $(x,y,j-1)$ and $(x,y,j+1)$, and the two which arise from the resulting data qubit error, at coordinates $(x,y,j)$ and $(x^{\prime},y^{\prime},j)$. This pattern is shown in Figure 7, together with its decomposition into two graph-like errors. Reference [28] contains an example decoding graph for a $4\times 4$ stability experiment using the error-spreading circuit that we generated with stim. In Figure 8, we show a $4\times 4$ planar code stability patch and the qubit pairs which take part in error-spreading events. In the bulk of the code, we require no additional qubits. However, on the boundary, we require $O(w)$ additional qubits for a $w\times w$ patch to ensure each auxiliary qubit has a qubit to which to spread its classification error; cf. Figure 1(b).
@@
We simulated the error-spreading circuits in order to compare the performance to standard syndrome extraction circuits with no reset. We applied the same principles for the number of shots as in Section II. We present the quantum memory results in Figure 9 and see that the performance is somewhat (although not significantly) diminished from the no-reset case. We attribute this to the error-spreading feature and the slightly higher depth of the error-spreading circuits. However, as we see from the stability time overhead plots Figure 10(b), we need fewer QEC rounds for the error-spreading circuit to avoid undetectable time-like logical failures and so we need fewer QEC rounds for quantum memory as well, making the gap in Figure 9 smaller. For instance, in case of the $4\times 4$ patch for $p=0.001$, we have $R_{\mathrm{spr}}\approx 0.85$, meaning we have to spend only $85\%$ of the time on stabiliser measurements as with the standard no-reset circuits. Taking into account the differing execution times, we see that we need only $\approx 80\%$ of the number of QEC rounds with the error-spreading circuit. Based on this, we can conclude that the error-spreading circuit improves the QEC performance over the standard no-reset circuits.
From comparing Figure 10(b) and Figure 5(c), we can further conclude that the error-spreading circuit is a good alternative to unconditional reset if the reset duration exceeds $100$ ns for the error regimes we considered. However, if the reset duration is decreased further, then unconditional reset outperforms the error-spreading circuit.
@@
Figure 10: Comparison of stability experiment performances in the no-reset scheme using standard (triangles and dashed lines) and error-spreading (squares and dash-dotted lines) circuits. Colours indicate different physical error probabilities: $p=10^{-2}$ (blue), $p=10^{-2.5}$ (orange), $p=10^{-3}$ (purple). (a) Comparison of logical failure probabilities for the $4\times 4$ stability patch. Error bars as in Figure 5(a) are also shown as shaded areas. Note that for the $p=10^{-2}$ case, the two lines overlap. (b) The time overhead $R_{\mathrm{spr}}$ that the error-spreading circuit requires when compared to the no-reset standard circuit.

### v2

@@ Abstract
Whether to reset qubits, or not, during quantum error correction experiments is a question of both foundational and practical importance for quantum computing. Text-book quantum error correction demands that qubits are reset after measurement. However, fast qubit reset has proven challenging to execute at high fidelity. Consequently, many cutting-edge quantum error correction experiments are opting for the no-reset approach, where physical reset is not performed. It has recently been postulated that no-reset is functionally equivalent to reset procedures, as well as being faster and easier. For memory experiments, we confirm numerically that resetting provides no benefit. On the other hand, we identify a remarkable difference during logical operations. We find that unconditionally resetting qubits can reduce the duration of fault-tolerant logical operation by up to a factor of two as the number of measurement errors that can be tolerated is doubled. We support this with numerical simulations. However, our simulations also reveal that the no-reset performance is superior if the reset duration and infidelity exceed given thresholds. For example, with the noise model we considered, we find the no-reset performance to be superior when the reset duration is greater than approximately $100$ ns and the physical error probability is greater than approximately $10^{-2.5}\approx 0.003$. Lastly, we introduce two novel syndrome extraction circuits that can reduce the time overhead of no-reset approaches. Our findings provide guidance on how experimentalists should design future experiments.
@@
The structure of our paper is as follows. In Section I, we describe how the decoding problem changes when we switch from the unconditional-reset to the conditional-reset or no-reset scheme. In Section II, we first describe a Pauli noise model, inspired by currently available realistic superconducting hardware, that we use throughout (Section II.1). Then, we present numerical results that verify that the differences between the three reset schemes for quantum memory with the rotated planar code (Section II.2) are small. After that, we describe the stability experiment and why it is a useful proxy for estimating the time cost of FTQC. Furthermore, we demonstrate through numerical simulations with the stability experiment that FTQC incurs a substantial time overhead in the no-reset (and conditional-reset) scheme, provided resetting is fast enough and is of high enough fidelity (Section II.3). In Sections III and IV, we present two alternative syndrome extraction circuits in the no-reset scheme that overcome this time overhead problem. Even though we present circuits only for the planar code, they can also be applied straightforwardly to more general QEC codes. With Section V, we conclude the paper by discussing the obtained results and possible future research directions.
@@ I The structure of detectors in the three reset schemes
Figure 4: Decoding graphs for measurements from a single auxiliary qubit in the unconditional-reset (left) and no-reset (right) cases.
Now, we analyse which detectors are triggered (change value from the noise-free case) by different errors. A Pauli error on a data qubit occurring during the initialisation of the auxiliary qubits triggers only detectors with the same time coordinate. We refer to errors with this property as “space-like”. On the other hand, a measurement error (due to either quantum bit flip or classical misclassification) on the auxiliary qubit of $g$ in the $j$th QEC round triggers at most two detectors, namely at coordinates $(x,y,j-1)$ and/or $(x,y,j)$, as shown on the left of Figure 4. Such an error is described as “time-like”. Some Pauli errors that occur mid-syndrome-extraction-circuit trigger a set of detectors where neither the spatial nor the time coordinates coincide for all the detectors – these are known as hook errors [24].
@@
This structural difference of detectors does not change the triggered detectors corresponding to any mid-circuit quantum Pauli error. However, misclassification of a measurement outcome results in a different combination of triggered detectors to the unconditional-reset case. To see this, let us assume we have a classification error on measurement result $n_{g,j}$, i.e., we read out $n_{g,j}\oplus 1$, even though the qubit collapsed into the $n_{g,j}$-eigenstate. This triggers two detectors at coordinates $(x,y,j-1)$ and $(x,y,j+1)$ (cf. Equation 2), so in the decoding graph it corresponds to a time-like edge with length $2$. In contrast, even in the no-reset case, a quantum measurement error corresponds to a time-like edge with length 1. A bit-flip on the auxiliary qubit associated with stabiliser $g$ just before measurement in the $k$th QEC round flips all measurements $n_{g,j}$ with $j\geq k$. This therefore triggers the detectors at coordinates $(x,y,k-1)$ and $(x,y,k)$. Both types of error are shown on the right of Figure 4.
In Figure 2, we present a topological perspective on how vertical strings of classification errors can lead to an undetectable logical failure in lattice surgery and stability experiments. In a lattice surgery operation, we measure a joint logical Pauli between 2 or more logical qubits, and the result is determined from a (corrected) product of stabiliser measurement outcomes. An uncorrected vertical string of classification errors will flip the value of one of these stabiliser measurements, and therefore flip the outcome of this logical Pauli measurement. The need to suppress such failure modes is why $d$ QEC rounds is the standard recommendation. When lattice surgery is used to perform a non-Clifford gate (e.g. a $T$ gate), we apply a logical Clifford gate conditional on the outcome of the lattice surgery measurement result. In such a situation, a logical measurement error during lattice surgery would be converted into a logical qubit error.
Such lattice surgery circuits are large and complex. Fortunately, we can instead use the smaller and simpler stability experiment, in simulation or on real qubits, to quantify the probability of logical failures due to a vertical string of errors. Stability experiments are therefore an excellent proxy for lattice surgery operations. Considering these vertical failure mechanisms, we see that the no-reset approach will tolerate only half as many classification errors as the unconditional-reset approach, since each error has twice the vertical length in the former scheme. This insight applies equally to lattice surgery and stability experiments. As we will see in Section II.3, this difference significantly impacts the performance of the stability experiment, and hence lattice-surgery-based FTQC. For other types of logical operation such as transversal gates, logical failure modes are less well understood – there is no known proxy in the same spirit as stability – and the required number of QEC rounds is an ongoing topic of debate [27, 28, 29] that we regard as unsettled. For quantum memory, the no-reset approach does not reduce the weight of minimal-weight logical errors, and so in this context we should expect only a minor variation in logical error rates.
@@ II.1 Noise model
We assume that the device has fixed and uniform qubit connectivity, with a planar nearest-neighbour architecture connecting each (bulk) qubit to four others. This so-called square-grid connectivity enables implementation of the planar code and is found in several modern superconducting quantum computers [e.g., 30, 14]. In Section IV, we will present a circuit that requires slightly different planar connectivity – some qubits need only be connected to three others instead of four.
@@
We construct our simulated circuits [31] in terms of the controlled-$Z$ two-qubit gate (CZ) and single-qubit $X$- and $Z$-basis rotations, a common native gate set for superconducting qubits [e.g., 30]. In particular, we use the $\pi/2$ rotations $\sqrt{X}$ and $S$, given by
@@
$T_{1}=\frac{p^{\mathrm{ref}}}{p}T_{1}^{\mathrm{ref}},\quad T_{2}=\frac{p^{\mathrm{ref}}}{p}T_{2}^{\mathrm{ref}}$ (6)
@@ II.2 The impact of no-reset on memory
Figure 5: Comparison of quantum memory performances in the unconditional-reset (dots and solid lines) and no-reset (triangles and dashed lines) schemes using standard syndrome extraction circuits. For each distance $d=5,7,9$ (corresponding to red, blue, black colours), we performed $d$ QEC rounds. The numbers of samples taken to calculate the data points are sufficiently large for the error bars not to be visible. As can be seen, there is only a small difference between the two schemes, confirming the statement made in [13, Sec. S2]. Note that the conditional reset scheme is not shown as the simulation is identical to the no-reset simulations except with higher noise levels due to the increased physical qubit idling. As a consequence, the conditional-reset scheme will always perform worse and is therefore not shown here.
Using the noise model described in Section II.1, we performed quantum memory simulations for distance-$d$ rotated planar code for $d$ number of QEC rounds. Figure 5 compares the unconditional-reset and no-reset schemes. We constructed circuits using the python library stim [34] and took samples which were then decoded using the MWPM python library pymatching [25]. The circuits are available here: [31]. The number of samples for each data point was the minimum required to observe $10^{4}$ logical failures or reach $10^{8}$ samples, whichever happened first. This ensures error bars that are too small to be visible in our plots. We calculated the sample logical failure probabilities $p_{X}$ and $p_{Z}$, for $X$ and $Z$-memory, respectively, and combined them into one quantity as $p_{L}=p_{X}+p_{Z}-p_{X}p_{Z}$. Our results (Figure 5) confirm the statement of [13], mentioned earlier, that there is only a small difference between the two reset schemes in the context of quantum memory. More precisely, the no-reset scheme’s performance is slightly better, as the threshold is somewhat higher and the logical failure probabilities are slightly lower. One reason for this difference is that the idle noise on the data qubits is lower in the no-reset case.
@@
Figure 6: Comparison of the stability experiment performances in the two reset schemes using standard syndrome extraction circuits. Colours indicate different physical error probabilities: $p=10^{-2}$ (blue), $p=10^{-2.5}$ (orange), $p=10^{-3}$ (purple). (a) Logical failure probabilities of the stability experiment for the $4\times 4$ patch shown in Figure 1(b) plotted against the number of QEC rounds. Error bars corresponding to 3×3$\times$3 × the standard error of the mean are also shown as shaded areas. Note that for the $p=10^{-2}$ case the plots overlap. The unconditional reset is plotted with dots, the no-reset with triangles. Best $\log$-line fits are also shown with solid and dashed lines, respectively. (b) The time overhead $R_{\mathrm{ur}}$ that the unconditional-reset scheme requires when compared to the no-reset scheme given $t_{\mathrm{res}}=0$ (instantaneous), $100$ (fast) and $500$ ns (slow) resets, respectively. In particular, the first column of dots in the first plot corresponds to (a), since instantaneous reset implies equal execution time in the two reset schemes. The shaded areas correspond to $90\%$ confidence intervals. The pink dashed line shows the break-even point, i.e. when $R_{\mathrm{ur}}=1$, below which the unconditional reset has better FTQC performance. Note that the conditional reset scheme is not shown as the simulation is identical to the no-reset simulation except with higher noise levels due to the increased physical qubit idling. As a consequence, the conditional reset scheme will always perform worse and is therefore not shown here.
@@ II.3 The impact of no-reset on FTQC
Based on this, we would expect the stability performance to be improved by using unconditional resets – assuming fast enough and high-enough fidelity reset gates. We performed simulations to assess this. We considered planar code stability patches of sizes $w\times w$ with $w\in\{4,6,8,10,12,14,16\}$ and physical error rates $p\in\{10^{-2},10^{-2.5},10^{-3}\}$. For each, we prepared stim circuits [31], sampled, and decoded until we reached either $10^{9}$ total shots or $10^{6}$ logical failures, whichever happened first. The obtained logical failure probabilities plotted against the number of QEC rounds, $n=5,7,9,11,13$, are shown in Figure 6(a) for the $w=4$ case. They clearly show that implementing unconditional reset improves the performance for $p=10^{-2.5},10^{-3}$. Furthermore, this improvement increases as $p$ decreases.
@@
where $a$ and $\gamma$ both depend on the width of the patch $w$ [21, Section 3]. The $\gamma$ parameter quantifies the strength of the exponential error suppression with the number of rounds in the stability experiment. Note that one QEC round takes $t_{\mathrm{nr}}=840$ ns with the no-reset scheme and, if the reset duration is $t_{\mathrm{res}}$ ns, one QEC round takes $t_{\mathrm{ur}}=(840+t_{\mathrm{res}})$ ns with the unconditional-reset scheme. These two times are equal only with instantaneous reset, i.e., $t_{\mathrm{res}}=0$ ns. Therefore, as a fairer comparison, Equation 7 may be expressed in terms of the time $t$ the experiment requires, instead of the number of QEC rounds, so that
@@
where subscripts $\mathrm{nr}$ and $\mathrm{ur}$ indicate the no- and unconditional-reset schemes, respectively. We define the time overhead of implementing unconditional resets as $R_{\mathrm{ur}}=\gamma_{\mathrm{nr}}/\gamma_{\mathrm{ur}}$ [35, Sec. 5.2]. This can be interpreted as follows: if we perform a stability experiment for $t_{\mathrm{nr}}$ ns without unconditional resets then, in order to match this performance using unconditional resets, we need to spend approximately $t_{\mathrm{ur}}=R_{\mathrm{ur}}t_{\mathrm{nr}}$ ns. If $R_{\mathrm{ur}}<1$, then implementing unconditional resets decreases the time cost of FTQC, and hence is worth implementing; otherwise not. The time overhead is plotted in Figure 6(b), assuming instantaneous reset $t_{\mathrm{res}}=0$ ns, fast reset $t_{\mathrm{res}}=100$ ns, and slow reset $t_{\mathrm{res}}=500$ ns. The plotted error bars correspond to $90\%$ confidence intervals. More precisely, for each data point, we sampled $1000$ times from a normal distribution with mean given by the sampled logical error probability, and standard deviation given by the standard error of the mean. Then, we took the best $\log$-line fits for each of the $1000$ cases, calculated the corresponding time overheads, and removed the $50$ smallest and the $50$ largest values hence obtained. The minimum and maximum of the remaining values are the bottom and top error bars, respectively. We call $p_{\mathrm{br}}$ the break-even point, where $R_{\mathrm{ur}}=1$. Clearly, as $t_{\mathrm{res}}$ increases, $p_{\mathrm{br}}$ decreases. Furthermore, it can be seen in Figure 6(b) that, as $p$ decreases, the time overhead decreases too, indicating that when $p$ is small enough, implementing unconditional reset improves FTQC performance. In all cases of instantaneous and fast reset, the break-even point satisfies $10^{-2.5}<p_{\mathrm{br}}<10^{-2}$. With slow reset, the break-even point is not visible in our plot, and we expect it to be at very low $p$. This indicates that, unless the reset gate error rate and duration are below certain values, implementing unconditional reset may not be beneficial for FTQC. However, if the break-even point is reached, either with decreased reset duration or higher-fidelity reset gates, the unconditional-reset scheme substantially improves FTQC performance.
@@ III Recovering time-like distance by spreading classification errors
The idea is to spread the auxiliary qubit measurement classification errors to the data qubits, thereby triggering additional detectors and thus requiring more errors for an undetectable time-like logical failure. We achieve this by applying two conditional Pauli gates immediately after measuring the auxiliary qubit – one conditioned on the measurement outcome and the other conditioned on the auxiliary qubit state itself; see Figure 7 for an $X$-stabiliser circuit. Note that the controlled gates are of $Z$-($X$-)type for $X$-($Z$-)type stabilisers, and that the classically-controlled gates do not have to be applied on the device – instead, their effect can be tracked in software. In the absence of a measurement classification error, the two conditional gates cancel so there is no overall effect on the data qubits. However, if there is a classification error, exactly one of the gates will have an effect, resulting in a Pauli gate being applied to the data qubit which then behaves like a Pauli error on the qubit. Therefore, misclassification of measurement $n_{g,j}$ will now trigger four detectors – the two which directly result from the measurement classification error, at coordinates $(x,y,j-1)$ and $(x,y,j+1)$, and the two which arise from the resulting data qubit error, at coordinates $(x,y,j)$ and $(x^{\prime},y^{\prime},j)$. This pattern is shown in Figure 8, together with its decomposition into two graph-like errors. Reference [31] contains an example decoding graph for a $4\times 4$ stability experiment using the error-spreading circuit that we generated with stim. In Figure 9, we show a $4\times 4$ planar code stability patch and the qubit pairs which take part in error-spreading events. In the bulk of the code, we require no additional qubits. However, on the boundary, we require $O(w)$ additional qubits for a $w\times w$ patch to ensure each auxiliary qubit has a qubit to which to spread its classification error; cf. Figure 1(b).
@@
We simulated the error-spreading circuits in order to compare the performance to standard syndrome extraction circuits with no reset. We applied the same principles for the number of shots as in Section II. We present the quantum memory results in Figure 10 and see that the performance is somewhat (although not significantly) diminished from the standard no-reset case. We attribute this to the error-spreading feature and the slightly higher depth of the error-spreading circuits. However, as we see from the stability time overhead plots Figure 11(b), we need fewer QEC rounds for the error-spreading circuit to avoid undetectable time-like logical failures and so we need fewer QEC rounds for quantum memory as well, making the gap in Figure 10 smaller. For instance, in case of the $4\times 4$ patch for $p=0.001$, we have $R_{\mathrm{spr}}\approx 0.85$, meaning we have to spend only $85\%$ of the time on stabiliser measurements as with the standard no-reset circuits. Taking into account the differing execution times, we see that we need only $\approx 80\%$ of the number of QEC rounds with the error-spreading circuit. Based on this, we can conclude that the error-spreading circuit improves the QEC performance over the standard no-reset circuits.
From comparing Figure 11(b) and Figure 6(b), we can further conclude that the error-spreading circuit is a good alternative to unconditional reset if the reset duration exceeds $100$ ns for the error regimes we considered. However, if the reset duration is decreased further, then unconditional reset outperforms the error-spreading circuit.
@@
Figure 11: Comparison of stability experiment performances in the no-reset scheme using standard (triangles and dashed lines) and error-spreading (squares and dash-dotted lines) circuits. Colours indicate different physical error probabilities: $p=10^{-2}$ (blue), $p=10^{-2.5}$ (orange), $p=10^{-3}$ (purple). (a) Comparison of logical failure probabilities for the $4\times 4$ stability patch. Error bars as in Figure 6(a) are also shown as shaded areas. Note that for the $p=10^{-2}$ case, the two lines overlap. (b) The time overhead $R_{\mathrm{spr}}$ that the error-spreading circuit requires when compared to the no-reset standard circuit. The shaded areas correspond to $90\%$ confidence intervals. The conditional reset scheme is not shown as the simulation is identical to the no-reset except with higher noise levels due to the increased physical qubit idling. As a consequence, the conditional reset scheme will always perform worse and is therefore not shown here.